OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. x + 9y = -26 -2x + 4y = -36

OpenStudy (anonymous):

do you mean the Gaus-Jordan Elimination method?

OpenStudy (anonymous):

not sure. It just says "Elimination method"

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

so its the other one which doesn't involve matrices

OpenStudy (anonymous):

I dont even know what that means :/

OpenStudy (anonymous):

 now you found the y value, we need to solve for x. You could use the 1st or 2nd equation for this but to make things simple, I'll go for the first one. x+9y=-36 x=-9y-36 x=-9(-4)-36 x=36-36 x=0